According to Yukawa theory of Nuclear forces, the attractive force between two Nucleons has the potential V(r)=Ke^-ar/r What will be the Nuclear force?


1
Expert's answer
2022-01-26T07:40:50-0500

"U(r)=\frac{Ke^{-ar}}r,"

"F=U'=\frac{-arKe^{-ar}-Ke^{-ar}}{r^2}=\frac{-Ke^{-ar}(1+ar)}{r^2}."